### v1

@@ 2.3 Strong isometric path complexity
A result of [6] implies that the strong isometric path complexity of graphs without a theta, pyramid, or a long prism is at most $71$. Now Proposition 2, imply the following:
@@
$r$$a_{1}=u$$b_{1}$$a_{2}$$b_{1}$$b_{2}$$a_{3}$$b_{3}=v$
Figure 1: Illustration of the proof of Lemma 7. Here $z=3$. The dashed curve from $u$ to $v$ indicates a $(u,v)$-path in $Y$. The solid black piecewise linear curve between $(u,v)$ that doesn’t pass through $r$, indicates the $(u,v)$-isometric path $P$ in $X$. For $p\in\{a_{1}=u,b_{2},b_{3}=v\}$, the strictly monotone solid black curves between $r,p$ indicate a $(r,p)$-isometric path in $X$.
@@
Let $P$ be any $(u,v)$-isometric path in $X$ and $\mathcal{Q}$ be a minimum cardinality set of $r$-rooted isometric paths in $X$ that covers $P$ and for any $Q\in\mathcal{Q}$, $|V(P)\cap V(Q)|$ is maximized. See Figure 1 for illustration. Let $z\coloneq|\mathcal{Q}|$ and by definition, $z\leq sipco\left(X\right)$. Observe that for any $Q\in\mathcal{Q}$, the path induced by $V(P)\cap V(Q)$ is an isometric path in $X$. Moreover, there exists exactly one path $Q_{1}\in\mathcal{Q}$ that contains $u$. Otherwise, if there exists a path $Q^{\prime}\in\mathcal{Q}\setminus\{Q_{1}\}$ containing $u$, then either $V(P)\cap V(Q^{\prime})$ contains $V(P)\cap V(Q_{1})$, or vice versa. But this contradicts the minimality of $\mathcal{Q}$.
@@
3 Proof of Theorem 3
For the entirety of this section, $G$ shall denote a fixed $K_{2,3}$-induced minor-free graph, and $r$ shall denote a fixed vertex of $G$. All clusters (and their parent sets) are defined w.r.t the metric in $G$ and w.r.t the vertex $r$. The output of Algorithm 1 with $G$ and $r$ as input will be denoted as $H$. See Figure 2 for an example. The sets $N_{k}(r)$, $B_{k}\left({r}\right)$, $P\left(r,S\right)$ always contains vertices of $G$ and are defined w.r.t the metric in $G$. In Section 3.1, we observe some properties of $H$. In Section 3.2, we prove lemmas relating the metric of $G$ and $H$. In Section 3.3, we prove some properties on the induced cycles of $H$, which we use in Section 3.4 to show that $H$ has tree-width at most two. We complete the proof in Section 3.5.
@@ 3 Proof of Theorem 3
Figure 2: The shaded regions indicate the clusters w.r.t $r$. The graph $H$ is the output of Algorithm 1 with $G$ and $r$ as input.
@@
2Let $E^{\prime}=\emptyset$.
3Construct $T=\Gamma\left({r},{G}\right)$.
4for each cluster $S\in V(T)$  do
@@ 3.1 Basic facts about $H$
Observe that, for every cluster $S$ of $G$ and vertex $v\in S$, we have that $v$ is adjacent to at least one vertex from $P\left(r,S\right)$ in $H$. Hence, the graph $H$ is connected. Below we state some observations which follow immediately from Algorithm 1. We provide the proofs for completeness.
@@
$(a)$ follows immediately from Algorithms 1 and 1 of Algorithm 1. To prove $(b)$, first consider the case when $F\coloneq G\left[P\left(r,S\right)\right]$ is connected. Clearly, $v$ has a neighbor $w\in P\left(r,S\right)$ such that $u$ and $w$ lie in the same connected component in $F$. Now consider the case when $F$ has two connected components $C_{1},C_{2}$ and for $i\in\{1,2\}$, let $D_{i}\coloneq N(C_{i})\cap S$. Since $C_{1},C_{2}$ are cliques in $G$ (due to Lemma 5(c)), and since there exists $i\in\{1,2\}$ such that $u\in C_{i}$, Algorithm 1 ensures that $v\in D_{i}$. This implies $C_{i}$ has a vertex $w$ with $\{vw,uw\}\subseteq E(G)$. ∎
1 immediately implies the following.
@@
The proof follows from Algorithms 1 and 1 of Algorithm 1. ∎
@@
The proof follows from Algorithm 1 of Algorithm 1. ∎
@@
The proof follows from Algorithm 1. ∎
@@ 3.2 Distances in $G$ and $H$
Let $k=d_{G}\left(r,x\right)=d_{G}\left(r,y\right)$, and $S\subseteq N_{k}(r)$ be the cluster containing $u,v$. Due to Lemma 5, the graph $F\coloneq G\left[P\left(r,S\right)\right]$ has at most two connected components. If $F$ is connected, then consider the vertex $w$ which is chosen in Algorithm 1 of Algorithm 1. Clearly, $\{xw,yw\}\subseteq E(H)$ (due to Algorithm 1 of Algorithm 1), and therefore $u,v$ has a common neighbor in $H$. Hence, $d_{H}\left(x,y\right)\leq 2$. Suppose $F$ has two connected components $C_{1},C_{2}$. For $i\in\{1,2\}$, let $D_{i}\coloneq\{u\in S\colon\exists v\in C_{i},uv\in E(G)\}$.
Suppose $D_{1}\cap D_{2}=\emptyset$. If there exists a $i\in\{1,2\}$, such that both $x,y\in D_{i}$, then there is a vertex $w\in C_{i}$ such that $\{xw,yw\}\subseteq E(H)$. Otherwise, due to Algorithm 1, there exists an edge $uv\in E(H)$ such that $u,x$ has a common neighbor in $H$ and $v,y$ has a common neighbor in $H$. Hence, $d_{H}\left(x,y\right)\leq 5$.
Now assume $D_{1}\cap D_{2}\neq\emptyset$. Without loss of generality assume that, the vertices of $C_{1}$ are colored as blue in Algorithm 1. If $\{x,y\}\subseteq D_{1}$ or $\{x,y\}\subseteq D_{2}\setminus D_{1}$, then $x,y$ has a common neighbor in $H$ (due to Algorithms 1 and 1 of Algorithm 1), and hence, $d_{H}\left(x,y\right)=2$. Otherwise, without loss of generality assume $x\in D_{1}$ and $y\in D_{2}\setminus D_{1}$. Due to Algorithm 1 of Algorithm 1, there exists a vertex $w\in D_{1}$ such that $x,w$ has a common neighbor in $H$ and $y,w$ has a common neighbor in $H$. Hence $d_{H}\left(x,y\right)\leq 4$. ∎
Next, we extend Lemma 9 and show that adjacent vertices in $G$ remain close to each other in $H$.
@@
Let $xy\in E(G)$, and without loss of generality assume that $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$. Let $k=d_{G}\left(r,x\right)$, and $S\subseteq N_{k}(r)$ be the cluster containing $x$. If $d_{G}\left(r,y\right)=k$, then $y\in S$ and we are done by Lemma 9. Now consider the case when $d_{G}\left(r,y\right)=k-1$. By definition, $y\in P\left(r,S\right)$. Due to Lemma 5, $F\coloneq G\left[P\left(r,S\right)\right]$ has at most two connected components. If $F$ is connected, then let $w$ be the vertex chosen in Algorithm 1 of Algorithm 1. Note that $xw\in E(H)$. Due to Lemma 5(b), there is a $(y,w)$-path $P\subseteq F$ of length at most three. Applying Lemma 9 on the each edge of the path $P$, we have that $d_{H}\left(x,y\right)\leq 16$.
Suppose $F$ has two connected components $C_{1},C_{2}$. Due to Lemma 5(c), each of $C_{1}$ and $C_{2}$ are cliques in $G$. For $i\in\{1,2\}$, let $D_{i}\coloneq\{u\in S\colon\exists v\in C_{i},uv\in E(G)\}$. Now we have the following cases.
If $D_{1}\cap D_{2}=\emptyset$, then there exists an integer $i\in\{1,2\}$ such that $x\in D_{i}$ and $y\in C_{i}$. In this case, there exists a vertex $w\in C_{i}$ such that $xw\in E(H)$ (due to Algorithms 1 and 1). Applying Lemma 9 on the edge $yw\in E(G)$, we have that $d_{H}\left(x,y\right)\leq 6$.
Consider the case, when $D_{1}\cap D_{2}\neq\emptyset$. Without loss of generality assume that the vertices of $C_{1}$ are colored blue in Algorithm 1. If $y\in C_{1},x\in D_{1}$ or if $y\in C_{2},x\in D_{2}\setminus D_{1}$, then using analogous arguments as before we have that $d_{H}\left(x,y\right)\leq 6$. Now suppose $y\in C_{2}$ and $x\in D_{1}\cap D_{2}$. Then, due to Algorithm 1 of Algorithm 1, there exist vertices $w_{1},v,w_{2}$ such that $\{xw_{1},w_{1}v,vw_{2}\}\subseteq E(H)$ and $w_{2}\in C_{2}$. Now applying Lemma 9 on the edge $yw_{2}\in E(G)$, we have that $d_{H}\left(x,y\right)\leq 8$. ∎
Note that the graph $H$ (i.e. the output of Algorithm 1) may not be a subgraph of $G$. In the following lemma, we show that the adjacent vertices in $H$ are also close to each other in $G$.
@@
Let $xy$ be any edge of $H$. Without loss of generality assume $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$. If $d_{G}\left(r,x\right)=d_{G}\left(r,y\right)$, then Algorithm 1 implies that $xy\in E(G)$. Hence, the lemma is true in this case. Now assume $d_{G}\left(r,y\right)<d_{G}\left(r,x\right)$. Algorithm 1 ensures that $d_{G}\left(r,y\right)=d_{G}\left(r,x\right)-1$. Let $S$ be the cluster that contains $x$ and let $F\coloneq G\left[P\left(r,S\right)\right]$. Due to 1(a), $y\in V(F)$. Due to 1(b), there exists a vertex in $z\in V(F)$ such that $xz\in E(G)$ and $y,z$ lie in the same connected component in $F$. If $F$ is connected, then due to Lemma 5(b) we have that $d_{G}\left(z,y\right)\leq 3$ and, thus $d_{G}\left(x,y\right)\leq 4$. Otherwise, due to Lemma 5(c), $F$ must have two connected components both of which are cliques. This implies $yz\in E(G)$ and thus $d_{G}\left(x,y\right)\leq 2$. ∎
@@
Let $Q$ be the $(x,y)$-subpath of $P$. Assume that the length of $Q$ is at least two. (Otherwise, we are done by Lemma 10.) Without loss of generality assume $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$ and $Q\coloneq x_{\ell}x_{\ell+1}\ldots,x_{k}$, with $x_{\ell}\coloneq y$ and $x_{k}\coloneq x$. The above definitions imply $d_{G}\left(x,y\right)=d_{G}\left(x_{k},x_{\ell}\right)=k-\ell$. Observe that any cluster contains at most one vertex of $Q$. For $i\in[\ell,k]$, let $S_{i}$ denote the cluster that contains $x_{i}$. Let $j\in[\ell,k]$ be the minimum integer that satisfy one of the following:
@@
If $k=j$, then the claim is trivially true. So we assume $j<k$. Let $x^{\prime}_{k}\coloneq x_{k}$ and for each $i\in[j,k)$, let $x^{\prime}_{i}\in S_{i}$ denote a vertex such that $x^{\prime}_{i}x^{\prime}_{i+1}\in E(H)$. Let $R\coloneq x^{\prime}_{k}x^{\prime}_{k-1}\ldots x^{\prime}_{j}$. If $G\left[P\left(r,S_{j+1}\right)\right]$ is connected, then due to Lemma 5(b), there is a $(x^{\prime}_{j},x_{j})$-induced path $Q^{\prime}\subseteq F$ of length at most three in $G$. Applying Lemma 9, on each edge of $Q^{\prime}$ we have that $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 15$ and hence $d_{H}\left(x_{k},x_{j}\right)\leq k-j+15$.
Now suppose $G\left[P\left(r,S_{j+1}\right)\right]$ has two connected components $C_{1},C_{2}$ such that the sets $D_{1}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{1},uv\in E(G)\}$ and $D_{2}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{2},uv\in E(G)\}$ have a common vertex. If there exists $i\in\{1,2\}$ with $x_{j},x^{\prime}_{j}\in C_{i}$, then applying Lemma 9 on the edge $x_{j}x^{\prime}_{j}\in E(G)$, we have $d_{G}\left(x^{\prime}_{j},x_{j}\right)\leq 5$. Since $D_{1}\cap D_{2}\neq\emptyset$, due to Algorithm 1, there exist $u\in S_{j+1}$ and $v\in S_{j}$ such that $\{x^{\prime}_{j}u,uv\}\subseteq E(H)$ and $vx_{j}\in E(G)$. Since $\{v,x_{j}\}\subseteq S_{j}$, Lemma 9 implies $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 7$. Hence, $d_{H}\left(x_{k},x_{j}\right)\leq k-j+7$. ∎
@@
For each $i\in[\ell,j-1]$, let the two connected components of $G\left[P\left(r,S_{i+1}\right)\right]$ are $C^{1}_{i}$ and $C^{2}_{i}$. Due to Proposition 1, both $C^{1}_{i}$ and $C^{2}_{i}$ are cliques in $G$. For each $t\in\{1,2\}$, let $D^{t}_{i}\coloneq\{u\in S_{i+1}\colon\exists v\in C^{t}_{i},uv\in E(G)\}$. By definition, $D^{1}_{i}\cap D^{2}_{i}=\emptyset$. Now without loss of generality assume that $x_{i+1}\in D^{1}_{i}$. Therefore, any vertex in $S_{i}$ which is adjacent to $x_{i+1}$ in $G$ must lie in $C^{1}_{i}$. Therefore, $x_{i}\in C^{1}_{i}$. Now due to Algorithm 1, $x^{\prime}_{i}\in C^{1}_{i}$. Since, $C^{1}_{i}$ is a clique, we have that $x_{i}x^{\prime}_{i}\in E(G)$. ∎
Let $R_{1}$ be a $(x_{k},x_{j})$-isometric path in $H$, define $R_{2}\coloneq x^{\prime}_{j}x^{\prime}_{j-1}\ldots x^{\prime}_{\ell}$ and $R_{3}$ be a $(x^{\prime}_{\ell},x_{\ell})$-isometric path in $H$. Due to 12.1, length of $R_{1}$ is at most $k-j+15$. Length of $R_{2}$ is at most $j-\ell$. Due to 12.2 and Lemma 9, length of $R_{3}$ is at most $5$. Hence, the length of the path $R_{1}\cup R_{2}\cup R_{3}$ is at most $k-\ell+20$. Hence $d_{H}\left(x,y\right)\leq d_{G}\left(x,y\right)+20$. ∎
In the next lemma, we show that if a pair of vertices lie in a $r$-rooted isometric path in $H$, then their distance in $G$ is decreased by an additive factor of at most five. The proof is similar to that of Lemma 12.
@@
Without loss of generality, assume that $d_{H}\left(r,y\right)\leq d_{H}\left(r,x\right)$. If $xy\in E(H)$, then we are done by Lemma 11. Otherwise, let $Q$ be the $(x,y)$-subpath of $P$. Let $k\coloneq d_{H}\left(r,x\right)$, :-$\ell\coloneq d_{H}\left(r,y\right)$, $x_{k}\coloneq x$, $x_{\ell}=y$ and $Q\coloneq x_{k}x_{k-1}\ldots x_{\ell}$. Observe that for $i\in[\ell,k]$, $d_{H}\left(r,x_{i}\right)=i$. Due to Lemma 8, $d_{G}\left(r,x_{i}\right)=i$. Therefore, any cluster contains at most one vertex of $Q$. For $i\in[\ell,k]$, let $S_{i}$ be the cluster that contains $x_{i}$. Due to 1(a), for $i\in[\ell+1,k]$, $x_{i-1}\in P\left(r,S_{i}\right)$. Let $j\in[\ell,k]$ be the minimum integer that satisfy the following:

### v2

@@ 2.3 Strong isometric path complexity
The “isometric path complexity” of a connected graph $G$, denoted by $ipco\left(G\right)$, is the minimum integer $k$ such that there exists a vertex $v\in V(G)$ that satisfy the following property: any isometric path of $G$ can be covered by $k$ many $v$-rooted isometric paths.
Definition 4 ([8, 10]).
@@
A result of [8] implies that the strong isometric path complexity of graphs without a theta, pyramid, or a long prism is at most $71$. Now Proposition˜2, imply the following:
@@
[…]
Figure 1: Illustration of the proof of Lemma˜7. Here $z=3$. The dashed curve from $u$ to $v$ indicates a $(u,v)$-path in $Y$. The solid black piecewise linear curve between $(u,v)$ that doesn’t pass through $r$, indicates the $(u,v)$-isometric path $P$ in $X$. For $p\in\{a_{1}=u,b_{2},b_{3}=v\}$, the strictly monotone solid black curves between $r,p$ indicate a $(r,p)$-isometric path in $X$.
@@
Let $P$ be any $(u,v)$-isometric path in $X$ and $\mathcal{Q}$ be a minimum cardinality set of $r$-rooted isometric paths in $X$ that covers $P$ and for any $Q\in\mathcal{Q}$, $|V(P)\cap V(Q)|$ is maximized. See Figure˜1 for illustration. Let $z\coloneq|\mathcal{Q}|$ and by definition, $z\leq sipco\left(X\right)$. Observe that for any $Q\in\mathcal{Q}$, the path induced by $V(P)\cap V(Q)$ is an isometric path in $X$. Moreover, there exists exactly one path $Q_{1}\in\mathcal{Q}$ that contains $u$. Otherwise, if there exists a path $Q^{\prime}\in\mathcal{Q}\setminus\{Q_{1}\}$ containing $u$, then either $V(P)\cap V(Q^{\prime})$ contains $V(P)\cap V(Q_{1})$, or vice versa. But this contradicts the minimality of $\mathcal{Q}$.
@@
3 Proof of Theorem˜3
For the entirety of this section, $G$ shall denote a fixed $K_{2,3}$-induced minor-free graph, and $r$ shall denote a fixed vertex of $G$. All clusters (and their parent sets) are defined w.r.t the metric in $G$ and w.r.t the vertex $r$. The output of Algorithm˜1 with $G$ and $r$ as input will be denoted as $H$. See Figure˜2 for an example. The sets $N_{k}(r)$, $B_{k}\left({r}\right)$, $P\left(r,S\right)$ always contains vertices of $G$ and are defined w.r.t the metric in $G$. In Section˜3.1, we observe some properties of $H$. In Section˜3.2, we prove lemmas relating the metric of $G$ and $H$. In Section˜3.3, we prove some properties on the induced cycles of $H$, which we use in Section˜3.4 to show that $H$ has tree-width at most two. We complete the proof in Section˜3.5.
@@ 3 Proof of Theorem˜3
Figure 2: The shaded regions indicate the clusters w.r.t $r$. The graph $H$ is the output of Algorithm˜1 with $G$ and $r$ as input.
@@
21mm
3
4Let $E^{\prime}=\emptyset$.
5Construct $T=\Gamma\left({r},{G}\right)$.
6for each cluster $S\in V(T)$ do
@@ 3.1 Basic facts about $H$
Observe that, for every cluster $S$ of $G$ and vertex $v\in S$, we have that $v$ is adjacent to at least one vertex from $P\left(r,S\right)$ in $H$. Hence, the graph $H$ is connected. Below we state some observations which follow immediately from Algorithm˜1. We provide the proofs for completeness.
@@
$(a)$ follows immediately from Algorithms˜1 and 1 of Algorithm˜1. To prove $(b)$, first consider the case when $F\coloneq G\left[P\left(r,S\right)\right]$ is connected. Clearly, $v$ has a neighbor $w\in P\left(r,S\right)$ such that $u$ and $w$ lie in the same connected component in $F$. Now consider the case when $F$ has two connected components $C_{1},C_{2}$ and for $i\in\{1,2\}$, let $D_{i}\coloneq N(C_{i})\cap S$. Since $C_{1},C_{2}$ are cliques in $G$ (due to Lemma˜5(c)), and since there exists $i\in\{1,2\}$ such that $u\in C_{i}$, Algorithm˜1 ensures that $v\in D_{i}$. This implies $C_{i}$ has a vertex $w$ with $\{vw,uw\}\subseteq E(G)$. ∎
˜1 immediately implies the following.
@@
The proof follows from Algorithms˜1 and 1 of Algorithm˜1. ∎
@@
The proof follows from Algorithm˜1 of Algorithm˜1. ∎
@@
The proof follows from Algorithm˜1. ∎
@@ 3.2 Distances in $G$ and $H$
Let $k=d_{G}\left(r,x\right)=d_{G}\left(r,y\right)$, and $S\subseteq N_{k}(r)$ be the cluster containing $u,v$. Due to Lemma˜5, the graph $F\coloneq G\left[P\left(r,S\right)\right]$ has at most two connected components. If $F$ is connected, then consider the vertex $w$ which is chosen in Algorithm˜1 of Algorithm˜1. Clearly, $\{xw,yw\}\subseteq E(H)$ (due to Algorithm˜1 of Algorithm˜1), and therefore $u,v$ has a common neighbor in $H$. Hence, $d_{H}\left(x,y\right)\leq 2$. Suppose $F$ has two connected components $C_{1},C_{2}$. For $i\in\{1,2\}$, let $D_{i}\coloneq\{u\in S\colon\exists v\in C_{i},uv\in E(G)\}$.
Suppose $D_{1}\cap D_{2}=\emptyset$. If there exists a $i\in\{1,2\}$, such that both $x,y\in D_{i}$, then there is a vertex $w\in C_{i}$ such that $\{xw,yw\}\subseteq E(H)$. Otherwise, due to Algorithm˜1, there exists an edge $uv\in E(H)$ such that $u,x$ has a common neighbor in $H$ and $v,y$ has a common neighbor in $H$. Hence, $d_{H}\left(x,y\right)\leq 5$.
Now assume $D_{1}\cap D_{2}\neq\emptyset$. Without loss of generality assume that, the vertices of $C_{1}$ are colored as blue in Algorithm˜1. If $\{x,y\}\subseteq D_{1}$ or $\{x,y\}\subseteq D_{2}\setminus D_{1}$, then $x,y$ has a common neighbor in $H$ (due to Algorithms˜1 and 1 of Algorithm˜1), and hence, $d_{H}\left(x,y\right)=2$. Otherwise, without loss of generality assume $x\in D_{1}$ and $y\in D_{2}\setminus D_{1}$. Due to Algorithm˜1 of Algorithm˜1, there exists a vertex $w\in D_{1}$ such that $x,w$ has a common neighbor in $H$ and $y,w$ has a common neighbor in $H$. Hence $d_{H}\left(x,y\right)\leq 4$. ∎
Next, we extend Lemma˜9 and show that adjacent vertices in $G$ remain close to each other in $H$.
@@
Let $xy\in E(G)$, and without loss of generality assume that $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$. Let $k=d_{G}\left(r,x\right)$, and $S\subseteq N_{k}(r)$ be the cluster containing $x$. If $d_{G}\left(r,y\right)=k$, then $y\in S$ and we are done by Lemma˜9. Now consider the case when $d_{G}\left(r,y\right)=k-1$. By definition, $y\in P\left(r,S\right)$. Due to Lemma˜5, $F\coloneq G\left[P\left(r,S\right)\right]$ has at most two connected components. If $F$ is connected, then let $w$ be the vertex chosen in Algorithm˜1 of Algorithm˜1. Note that $xw\in E(H)$. Due to Lemma˜5(b), there is a $(y,w)$-path $P\subseteq F$ of length at most three. Applying Lemma˜9 on the each edge of the path $P$, we have that $d_{H}\left(x,y\right)\leq 16$.
Suppose $F$ has two connected components $C_{1},C_{2}$. Due to Lemma˜5(c), each of $C_{1}$ and $C_{2}$ are cliques in $G$. For $i\in\{1,2\}$, let $D_{i}\coloneq\{u\in S\colon\exists v\in C_{i},uv\in E(G)\}$. Now we have the following cases.
If $D_{1}\cap D_{2}=\emptyset$, then there exists an integer $i\in\{1,2\}$ such that $x\in D_{i}$ and $y\in C_{i}$. In this case, there exists a vertex $w\in C_{i}$ such that $xw\in E(H)$ (due to Algorithms˜1 and 1). Applying Lemma˜9 on the edge $yw\in E(G)$, we have that $d_{H}\left(x,y\right)\leq 6$.
Consider the case, when $D_{1}\cap D_{2}\neq\emptyset$. Without loss of generality assume that the vertices of $C_{1}$ are colored blue in Algorithm˜1. If $y\in C_{1},x\in D_{1}$ or if $y\in C_{2},x\in D_{2}\setminus D_{1}$, then using analogous arguments as before we have that $d_{H}\left(x,y\right)\leq 6$. Now suppose $y\in C_{2}$ and $x\in D_{1}\cap D_{2}$. Then, due to Algorithm˜1 of Algorithm˜1, there exist vertices $w_{1},v,w_{2}$ such that $\{xw_{1},w_{1}v,vw_{2}\}\subseteq E(H)$ and $w_{2}\in C_{2}$. Now applying Lemma˜9 on the edge $yw_{2}\in E(G)$, we have that $d_{H}\left(x,y\right)\leq 8$. ∎
Note that the graph $H$ (i.e. the output of Algorithm˜1) may not be a subgraph of $G$. In the following lemma, we show that the adjacent vertices in $H$ are also close to each other in $G$.
@@
Let $xy$ be any edge of $H$. Without loss of generality assume $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$. If $d_{G}\left(r,x\right)=d_{G}\left(r,y\right)$, then Algorithm˜1 implies that $xy\in E(G)$. Hence, the lemma is true in this case. Now assume $d_{G}\left(r,y\right)<d_{G}\left(r,x\right)$. Algorithm˜1 ensures that $d_{G}\left(r,y\right)=d_{G}\left(r,x\right)-1$. Let $S$ be the cluster that contains $x$ and let $F\coloneq G\left[P\left(r,S\right)\right]$. Due to ˜1(a), $y\in V(F)$. Due to ˜1(b), there exists a vertex in $z\in V(F)$ such that $xz\in E(G)$ and $y,z$ lie in the same connected component in $F$. If $F$ is connected, then due to Lemma˜5(b) we have that $d_{G}\left(z,y\right)\leq 3$ and, thus $d_{G}\left(x,y\right)\leq 4$. Otherwise, due to Lemma˜5(c), $F$ must have two connected components both of which are cliques. This implies $yz\in E(G)$ and thus $d_{G}\left(x,y\right)\leq 2$. ∎
@@
Let $Q$ be the $(x,y)$-subpath of $P$. Assume that the length of $Q$ is at least two. (Otherwise, we are done by Lemma˜10.) Without loss of generality assume $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$ and $Q\coloneq x_{\ell}x_{\ell+1}\ldots,x_{k}$, with $x_{\ell}\coloneq y$ and $x_{k}\coloneq x$. The above definitions imply $d_{G}\left(x,y\right)=d_{G}\left(x_{k},x_{\ell}\right)=k$-$\ell$. Observe that any cluster contains at most one vertex of $Q$. For $i\in[\ell,k]$, let $S_{i}$ denote the cluster that contains $x_{i}$. Let $j\in[\ell,k]$ be the minimum integer that satisfy one of the following:
@@
If $k=j$, then the claim is trivially true. So we assume $j<k$. Let $x^{\prime}_{k}\coloneq x_{k}$ and for each $i\in[j,k)$, let $x^{\prime}_{i}\in S_{i}$ denote a vertex such that $x^{\prime}_{i}x^{\prime}_{i+1}\in E(H)$. Let $R\coloneq x^{\prime}_{k}x^{\prime}_{k-1}\ldots x^{\prime}_{j}$. If $G\left[P\left(r,S_{j+1}\right)\right]$ is connected, then due to Lemma˜5(b), there is a $(x^{\prime}_{j},x_{j})$-induced path $Q^{\prime}\subseteq F$ of length at most three in $G$. Applying Lemma˜9, on each edge of $Q^{\prime}$ we have that $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 15$ and hence $d_{H}\left(x_{k},x_{j}\right)\leq k-j+15$.
Now suppose $G\left[P\left(r,S_{j+1}\right)\right]$ has two connected components $C_{1},C_{2}$ such that the sets $D_{1}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{1},uv\in E(G)\}$ and $D_{2}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{2},uv\in E(G)\}$ have a common vertex. If there exists $i\in\{1,2\}$ with $x_{j},x^{\prime}_{j}\in C_{i}$, then applying Lemma˜9 on the edge $x_{j}x^{\prime}_{j}\in E(G)$, we have $d_{G}\left(x^{\prime}_{j},x_{j}\right)\leq 5$. Since $D_{1}\cap D_{2}\neq\emptyset$, due to Algorithm˜1, there exist $u\in S_{j+1}$ and $v\in S_{j}$ such that $\{x^{\prime}_{j}u,uv\}\subseteq E(H)$ and $vx_{j}\in E(G)$. Since $\{v,x_{j}\}\subseteq S_{j}$, Lemma˜9 implies $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 7$. Hence, $d_{H}\left(x_{k},x_{j}\right)\leq k-j+7$. ∎
@@
For each $i\in[\ell,j-1]$, let the two connected components of $G\left[P\left(r,S_{i+1}\right)\right]$ are $C^{1}_{i}$ and $C^{2}_{i}$. Due to Proposition˜1, both $C^{1}_{i}$ and $C^{2}_{i}$ are cliques in $G$. For each $t\in\{1,2\}$, let $D^{t}_{i}\coloneq\{u\in S_{i+1}\colon\exists v\in C^{t}_{i},uv\in E(G)\}$. By definition, $D^{1}_{i}\cap D^{2}_{i}=\emptyset$. Now without loss of generality assume that $x_{i+1}\in D^{1}_{i}$. Therefore, any vertex in $S_{i}$ which is adjacent to $x_{i+1}$ in $G$ must lie in $C^{1}_{i}$. Therefore, $x_{i}\in C^{1}_{i}$. Now due to Algorithm˜1, $x^{\prime}_{i}\in C^{1}_{i}$. Since, $C^{1}_{i}$ is a clique, we have that $x_{i}x^{\prime}_{i}\in E(G)$. ∎
Let $R_{1}$ be a $(x_{k},x_{j})$-isometric path in $H$, define $R_{2}\coloneq x^{\prime}_{j}x^{\prime}_{j-1}\ldots x^{\prime}_{\ell}$ and $R_{3}$ be a $(x^{\prime}_{\ell},x_{\ell})$-isometric path in $H$. Due to ˜12.1, length of $R_{1}$ is at most kj+15k-j+15. Length of $R_{2}$ is at most $j$-$\ell$. Due to ˜12.2 and Lemma˜9, length of $R_{3}$ is at most $5$. Hence, the length of the path $R_{1}\cup R_{2}\cup R_{3}$ is at most $k-\ell+20$. Hence $d_{H}\left(x,y\right)\leq d_{G}\left(x,y\right)+20$. ∎
In the next lemma, we show that if a pair of vertices lie in a $r$-rooted isometric path in $H$, then their distance in $G$ is decreased by an additive factor of at most five. The proof is similar to that of Lemma˜12.
@@
Without loss of generality, assume that $d_{H}\left(r,y\right)\leq d_{H}\left(r,x\right)$. If $xy\in E(H)$, then we are done by Lemma˜11. Otherwise, let $Q$ be the $(x,y)$-subpath of $P$. Let $k\coloneq d_{H}\left(r,x\right)$, $\ell\coloneq d_{H}\left(r,y\right)$, $x_{k}\coloneq x$, $x_{\ell}=y$ and $Q\coloneq x_{k}x_{k-1}\ldots x_{\ell}$. Observe that for $i\in[\ell,k]$, $d_{H}\left(r,x_{i}\right)=i$. Due to Lemma˜8, $d_{G}\left(r,x_{i}\right)=i$. Therefore, any cluster contains at most one vertex of $Q$. For $i\in[\ell,k]$, let $S_{i}$ be the cluster that contains $x_{i}$. Due to ˜1(a), for $i\in[\ell+1,k]$, $x_{i-1}\in P\left(r,S_{i}\right)$. Let $j\in[\ell,k]$ be the minimum integer that satisfy the following: